a 6.00% coupon bonds have 9 years left to maturity. the bond makes annual coupon payments. if the yield to maturity on these bonds is

A 6.00% coupon bonds have 9 years left to maturity. The bonds make annual coupon payments . If the yield to maturity on these bonds is 8.00%, what is the current bond price
Answer #1

Annual coupon=1000*6%=60

Hence current price=Annual coupon*Present value of annuity factor(8%,9)+1000*Present value of discounting factor(8%,9)

=60*6.24688791+1000*0.500248967

=$875.06(Approx)

NOTE:

1.Present value of annuity=Annuity[1-(1+interest rate)^-time period]/rate

=60*[1-(1.08)^-9]/0.08

=60*6.24688791

2.Present value of discounting factor=1000/1.08^9

=1000*0.500248967